OpenStudy (anonymous):

Given the equation y − 3 = one half(x + 6) in point-slope form, identify the equation of the same line in standard form. x − 2y = −12 y = one halfx y = one halfx + 6 y = one halfx + 9

OpenStudy (anonymous):

@AravindG Help?

OpenStudy (aravindg):

just use distributive property on RHS and take all terms to LHS

OpenStudy (anonymous):

Huh?

OpenStudy (aravindg):

huh?

OpenStudy (aravindg):

just open bracket in Right hand side

OpenStudy (anonymous):

Will you show me how to do it im so confused.

OpenStudy (aravindg):

is ur equation \[y-3=\dfrac{1 }{2} \times (x+6) ?\]

OpenStudy (anonymous):

\[y - 3 = \frac{ 1 }{ 2 }(x + 6)\]

OpenStudy (aravindg):

ok open brackets on RHS what do you get?

OpenStudy (anonymous):

7?

OpenStudy (aravindg):

7?

OpenStudy (anonymous):

Oh wait, 3?

OpenStudy (aravindg):

yep \[y-3=\dfrac{x}{2}+3\]

OpenStudy (anonymous):

Now what

OpenStudy (aravindg):

now take all terms to LHS

OpenStudy (aravindg):

OR WAIT

OpenStudy (anonymous):

how?

OpenStudy (aravindg):

a better idea would be to add 3 on both sides

OpenStudy (aravindg):

\[y-3=\dfrac{x}{2}+3 \]

OpenStudy (aravindg):

adding 3 on both sides what do you get ?

OpenStudy (aravindg):

@ahoward79 ?

OpenStudy (anonymous):

y = x/2 + 6

OpenStudy (aravindg):

check the options now

OpenStudy (aravindg):

:)

OpenStudy (aravindg):

@ahoward79 got it?

OpenStudy (anonymous):

Oh! Yes! That was pretty easy :) Thank yu! Can you help me with one more?

OpenStudy (aravindg):

lols will try :D

OpenStudy (anonymous):

Since it wasn't corect can we look at it again?